# Teuken-7B-Base & Teuken-7B-Instruct: Towards European LLMs

### v2

@@ 7 Results
Section 7.1 describes our evaluation set-up, Section 7.2 presents our results across all 21 investigated languages, while Section 7.3 focuses on the performance of our models in the six widely spoken languages: English, German, French, Italian, Spanish, and Portuguese, and Section 7.4 presents our findings on the remaining 15 languages. Additionally, Section 7.5 highlights the performance of various multilingual models on the largest common language subset. Finally, Section 7.6 presents the instruction following capabilities of Teuken-7B-Instruct on the recently published multilingual benchmark MT-Bench-X Weber et al. (2024).
@@ 7.1 Evaluation Set-Up
The evaluation was conducted using ARC Clark et al. (2018)/EU21-ARC Thellmann et al. (2024) (25-shot for science-based questions), HellaSwag Zellers et al. (2019)/EU21-HeSw Thellmann et al. (2024) (10-shot for commonsense reasoning), TruthfullQA Lin et al. (2022)/EU21-TQA Thellmann et al. (2024) (6-shot pseudo-shot for generating truthful answers), and MMLU Hendrycks et al. (2021)/EU21-MMLU Thellmann et al. (2024) (5-shot for broad knowledge), which are available in 21 of the 24 official European languages, for which we report mean accuracy.
@@
We evaluated our models against related multilingual models with up to 8B parameters that have been pre-trained on causal language modelling from scratch and models that result from an instruction/fine-tuning of the pre-trained model. We did not include models that represent the distillation of larger models to ensure a comparable setting. As a result, we evaluated against Aya-23 8B, Bloom-7B1, Bloomz-7B1, Meta-Llama-3.1-8B, Meta-Llama-3.1-8B Chat, Salamandra-7B, Salamandra-7B Instruct, Pharia-1-LLM-7B, Pharia-1-LLM-7B-Control-Aligned, Occiglot-7B-EU5, and Mistral. Table 1 provides an overview of the number of languages and tokens that the models have been trained on. Note that the Mistral models, compared to those in this study, were not trained as multilingual models and are intended to be used as English-only models.
@@ 7.2 Performance on 21 European Languages
Table 2 and Table 9 present the multilingual results of the instruction-tuned and based models for all 21 European languages. Table 2 illustrates that Meta-Llama-3.1-8B-Instruct leads with an average accuracy of 56.3% across all benchmarks, followed by our instruction-tuned model and the Salamandra model. The results are encouraging, considering that our model has been trained on significantly fewer tokens than both other models (see Tables 1). We hypothesize that the training on the FineWeb-EDU dataset has contributed significantly to this aspect since it has shown to provide similar results compared to different composed datasets while requiring significantly fewer tokens Penedo et al. (2024). Noteworthy, it improves cross-lingual performance even though the dataset is in English (see Figure 4). On benchmark-level, Meta-Llama-3.1-8B-Instruct excels in EU21-MMLU (57.6%) while Salamandra-7B-Instruct is particularly strong in EU21-ARC (58.9%) and EU21-HeSw (63.7%) and Mistral-7B-Instruct-v0.3 obtains the best result on EU21-TQA (54.8%). Again, our model is competitive with the best-performing model in all tasks except for EU21-MMLU despite being trained on significantly fewer tokens.
Besides the average performance across languages, we also investigated the robustness across languages. Figure 5 shows that our model and Salamandra-7B-Instruct and Bloomz-7B1 are significantly more robust than the other models on all benchmarks except EU21-TQA where several models obtain robust performance across languages, ensuring that the model performs comparably across languages.
Model Average EU21-ARC EU21-HeSw EU21-TQA EU21-MMLU
@@
Ours (Base) .496 .550 .615 .469 .349
Ours (Instruct) .543 .581 .624 .543 .425
Table 2: Results on multilingual benchmarks for 21 European languages with instruction tuned models.
7.3 Performance on Top-6 European Languages
Table 3 focuses on the performance of instruction tuned models on six widely spoken European languages: English, German, French, Italian, Spanish, and Portuguese. These languages are well-represented in training data and are commonly used in evaluating multilingual models. In the instruction-tuned comparison, Meta-Llama-3.1-8B-Instruct leads with 62.3% average performance across tasks, excelling in EU21-HeSw (67.7%) and EU21-ARC (64.8%), followed closely by Mistral-7B-Instruct-v0.3 with 61.3%, obtaining the best performance in EU21-TQA (56.8%) and EU21-ARC (65.4%). Occiglot-7B-eu5-Instruct which is a fine-tuned Mistral model outperforms the all models on EU21-HeSw followed by Pharia-1-LLM-7B-control-aligned. Ours (Instruct) performs competitively in EU21-HeSw 67.5% and EU21-ARC (63.2%), but shows limitations in EU21-MMLU (45.7%), particularly in domain-specific tasks.
Model Average EU21-ARC EU21-HeSw EU21-TQA EU21-MMLU
@@ 7.3 Performance on Top-6 European Languages
Ours (Base) .521 .599 .667 .445 .371
Ours (Instruct) .574 .632 .675 .532 .457
Table 3: Results on multilingual benchmarks for 6 Languages (English, German, French, Italian, Spanish, Portuguese) with instruction tuned models.
7.4 Performance on Exclusive European Languages
Table 4 compares the models across 15 less commonly evaluated European languages (Romanian, Czech, Danish, Greek, Estonian, Finnish, Hungarian, Lithuanian, Latvian, Dutch, Bulgarian, Polish, Slovak, Slovenian, and Swedish). Meta-Llama-3.1-8B-Instruct leads with an average accuracy of 54.8%, excelling in EU21-MMLU (56.2%) and EU21-HeSw (55.6%). Ours (Instruct) closely follows with 53.6%, outperforming Meta-Llama in EU21-HeSw (61.3%) and EU21-TQA (54.5%), highlighting its strengths in commonsense reasoning and truthfulness. Salamandra-7B-Instruct stands out in EU21-ARC (58.5%) and EU21-HeSw (62.8%) with an average of 54.0%.
@@ 7.4 Performance on Exclusive European Languages
Meta-Llama-3.1-8B-Instruct .548 .542 .556 .531 .562
Mistral-7B-Instruct-v0.3 .505 .497 .508 .543 .473
Salamandra-7B-Instruct .540 .585 .628 .491 .455
Aya-23-8B .460 .435 .493 .477 .436
Occiglot-7B-eu5-Instruct .444 .437 .463 .474 .402
Pharia-1-LLM-7B-Aligned .369 .322 .362 .460 .331
Bloomz-7B1 .339 .281 .314 .473 .287
Ours (Base) .491 .540 .603 .474 .345
Ours (Instruct) .536 .569 .613 .545 .418
Table 4: Results on multilingual benchmarks across the 15 exclusive European languages (Romanian, Czech, Danish, Greek, Estonian, Finnish, Hungarian, Lithuanian, Latvian, Dutch, Bulgarian, Polish, Slovak, Slovenian, and Swedish.
@@ 7.5 Comparison on Common Languages
The tables  Table 12 and  Table 8 compare the performance of “Ours (Base)” and “Ours (Instruct)” models with Aya, Bloom, and Salamandra across ten common European languages: Czech, Dutch, English, French, Greek, Italian, Polish, Portuguese, Romanian, and Spanish. In the base model comparison, Ours (Base) performs well with an average accuracy of 51.2%, outperforming Bloom-7B1 (37.6%) across all tasks. Ours (Base) shines in EU21-HeSw with 64.3%, coming close to Salamandra-7B’s leading 66.4%, while also showing solid results in EU21-ARC (57.8%) and EU21-TQA (46.6%). In the instruct model comparison, Ours (Instruct) achieves 56.0%, closely matching Aya-23-8B’s top score of 56.3%. Ours leads in EU21-TQA (54.1%), outperforming Aya (48.4%) and Salamandra (47.6%), and remains competitive in EU21-ARC (61.0%) and EU21-HeSw (65.3%), making it a strong contender overall.
@@ 7.6 Multilingual Instruction Following Evaluation
For evaluating the multilingual instruction-following capabilities of Teuken-7B-Instruct, we utilized MT-Bench-X Weber et al. (2024) across all five available evaluation languages: English, German, French, Italian, and Spanish. The results are presented in the Table 41. We compare the cross-lingual performance of our model with Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3 and Salamandra-7B-Instruct in the following. While the creative writing skills (Writing, Roleplay) and question answering skills (Stem, Humanities) of our model show good performance and are comparable for German to Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, the mathematical, coding, and reasoning skills are inferior across languages. However, this shows to be a common pattern across all evaluated models. Compared to Salamandra-7B-Instruct, our model demonstrates superior cross-lingual capabilities, as shown in Figure 14(a). It is only at a language- and skill-specific level — namely in EN-Extraction and FR- and ES-Reasoning — that it outperforms our model.
@@ 8 Conclusion & Future Work
In this work, we presented the development of two multilingual large language models, “Ours (Base)” and “Ours (Instruct)”, tailored to support the linguistic diversity of Europe by encompassing all 24 official EU languages. Through the use of a custom multilingual tokenizer and a dataset prioritizing non-English content, our models address limitations found in existing multilingual models, particularly their English-centric bias. Our results demonstrate competitive performance across multiple benchmarks, including ARC, HellaSwag, MMLU, and TruthfulQA, while training on significantly fewer tokens. The presented multilingual models, focusing on European languages, are an excellent starting point for further pre-training or fine-tuning each European language. We have taken great care to ensure that all languages are represented in the tokenizer, promoting the inclusivity and applicability of our models in diverse linguistic contexts. We shared insights regarding our model development, often not described in detail, to support future development aimed at developing multilingual models.
Our ongoing efforts aim to improve the models’ capabilities further, focusing on math and coding capabilities by further training/fine-tuning our models. We emphasized the competitiveness of our model’s results despite training on fewer tokens. Therefore, we will focus on further improving the training datasets’ data quality to ensure efficient training, resulting in decreased resource requirements. Finally, we aim to train and share larger LLMs covering an even larger set of European languages beyond the 24 official European languages, making the technology available to an even larger community.
9 Limitations
Developing LLMs and particularly multilingual LLMs is a complex undertaking which involves several trade-offs during the different development phases (data composition, tokenize development, pre-training, instruction-tuning, and evaluation) Despite the essential contribution of our work towards European LLMs, it has several limitations.
First, we did not investigate multilingual scaling laws due to limited resources. This should be conducted in a separate effort, and our results present an excellent reference.
Second, our model has also been trained on educational content (FineWeb-EDU) available in English. Our results revealed that training on educational content in English improves the cross-lingual performance of the model. However, the performance varies across languages. We assume that employing educational content for other (low-resource) languages will strongly benefit these languages.
Third, we decided not to particularly focus on coding and math content, i.e., our model should not be employed in a context where these capabilities are required even though it has these capabilities to an extent.
Fourth, our model has a context window of 4096 tokens, currently limiting the use of the model in settings where long documents should be processed.
Fifth, due to the limited resources, our model has been trained on significantly fewer tokens than related models, which might explain why our model performs poorly in specific tasks.
Because we open source our models, the limitations 2-5 can be addressed by further training/instruction tuning the models dependent on the specific needs. Since our model relies on a multilingual tokenizer that ensures low fertility across languages, further training can be conducted efficiently.
10 Ethical & Broader Impact
Considering the enormous potential LLMs to be applied in various use cases, limiting them to a small number of languages introduces severe disadvantages such as lower performance, increased continued pre-training/fine-tuning, inference costs for various communities to relying on the selected languages. Our work makes a significant step towards democratizing this technology within Europe, which represents a diverse linguistic context. Furthermore, our strong focus on multilingualism encompasses the different cultural aspects reflected in natural language.
Because training LLMs requires vast resources, it is crucial to use them responsibly. By employing a multilingual tokenizer, multilingual documents are not unnecessarily highly fragmented, ensuring efficient training and reduced carbon footprint. Consequently, future work further training our models can be conducted efficiently across the investigated languages.

### v3

@@ 7 Results
Section˜7.1 describes our evaluation set-up, Section˜7.2 presents our results across all 21 investigated languages, while Section˜7.3 focuses on the performance of our models in the six widely spoken languages: English, German, French, Italian, Spanish, and Portuguese, and Section˜7.4 presents our findings on the remaining 15 languages. Additionally, Section˜7.5 highlights the performance of various multilingual models on the largest common language subset. Further results comparing our base and instruction-tuned model to related models across different language sets and individual languages are presented in the appendix (Table 7–Table 31). Finally, Section˜7.6 presents the instruction following capabilities of Teuken-7B-Instruct on the recently published multilingual benchmark MT-Bench-X [1].
@@ 7.1 Evaluation Set-Up
The evaluation was conducted using ARC [40]/EU21-ARC [67] (25-shot for science-based questions), HellaSwag [44]/EU21-HeSw [67] (10-shot for commonsense reasoning), TruthfullQA [47]/EU21-TQA [67] (6-shot pseudo-shot for generating truthful answers), and MMLU [16]/EU21-MMLU [67] (5-shot for broad knowledge), which are available in 21 of the 24 official European languages, for which we report mean accuracy.
@@
We evaluated our models against related multilingual models with up to 8B parameters that have been pre-trained on causal language modelling from scratch and models that result from an instruction/fine-tuning of the pre-trained model. We did not include models that represent the distillation of larger models to ensure a comparable setting. As a result, we evaluated against Aya-23 8B, Bloom-7B1, Bloomz-7B1, Meta-Llama-3.1-8B, Meta-Llama-3.1-8B Chat, Salamandra-7B, Salamandra-7B Instruct, Pharia-1-LLM-7B, Pharia-1-LLM-7B-C-A (Control-Aligned), Occiglot-7B-EU5, and Mistral. Table˜1 provides an overview of the number of languages and tokens that the models have been trained on. Note that the Mistral models were not trained as multilingual models and are intended to be used as English-only models. However, they contain multilingual to some extent, as described in [9]. Due to its strong performance, we included Mistral in our evaluations.
@@ 7.2 Performance on 21 European Languages
Table˜2(a) and Table˜7 present the multilingual results of the instruction-tuned and based models for all 21 European languages. Table˜2(a) illustrates that our instruction-tuned model leads with an average accuracy of 57.0% across all benchmarks, followed by Meta-Llama-3.1-8B-Instruct. The results are remarkable, considering our model has been trained on significantly fewer tokens than Llama-3.1-8B-Instruct (see Table 1). We hypothesize that the training on the FineWeb-EDU dataset has contributed significantly to this aspect since it has shown to provide similar results compared to different composed datasets while requiring significantly fewer tokens [20]. Noteworthy, it improves cross-lingual performance (see Figure˜5) even though the dataset is in English. On benchmark-level, Meta-Llama-3.1-8B-Instruct excels in EU21-MMLU (57.6%) while Salamandra-7B-Instruct is particularly strong in EU21-ARC (59.5%).
Besides the average performance across languages, we also investigated the robustness across languages. Figure˜6 shows that our model and Salamandra-7B-Instruct and Bloomz-7B1 are significantly more robust than the other models on all benchmarks except for EU21-TQA where several models obtain robust performance across languages, ensuring that the model performs comparably across languages.
7.3 Performance on Top-6 European Languages
Table˜2(b) focuses on the performance of the models on six widely spoken European languages: English, German, French, Italian, Spanish, and Portuguese. These languages are well-represented in training data and commonly used to evaluate multilingual models. Meta-Llama-3.1-8B-Instruct leads with 62.3% average performance across tasks, excelling in EU21-MMLU (63.2%), followed closely by Mistral-7B-Instruct-v0.3 with 61.3% on average, obtaining the best performance in EU21-TQA (56.8%) and EU21-ARC (65.4%). Teuken-7B-Instruct (Ours) outperforms all models on EU21-HeSw with 71.9% and performs competitively in all remaining benchmarks except for EU21-MMLU.
7.4 Performance on Exclusive European Languages
Table˜2(c) compares the models across 15 less commonly evaluated European languages (Romanian, Czech, Danish, Greek, Estonian, Finnish, Hungarian, Lithuanian, Latvian, Dutch, Bulgarian, Polish, Slovak, Slovenian, and Swedish). Teuken-7B-Instruct (Ours) leads with an average accuracy of 55.9%, excelling in EU21-HeSw (64.0%) and EU21-TQA (58.3%), followed by Meta-Llama-3.1-8B-Instruct that obtained the best performance on EU21-MMLU (55.4%) and Salamandra-7B-Instruct that obtained the strongest performance on on EU21-ARC (57.6%).
Model Avg. EU21-ARC EU21-HeSw EU21-TQA EU21-MMLU
@@ 7.4 Performance on Exclusive European Languages
Teuken-7B-Base (Ours) .520 .558 .619 .449 .453
Teuken-7B-Instruct (Ours) .570 .590 .663 .573 .454
(a) Results on multilingual benchmarks for 21 European languages with instruction-tuned models.
Model Avg. EU21-ARC EU21-HeSw EU21-TQA EU21-MMLU
@@
Teuken-7B-Base (Ours) .541 .608 .674 .405 .478
Teuken-7B-Instruct (Ours) .598 .641 .719 .549 .482
(b) Results on multilingual benchmarks for 6 Languages (English, German, French, Italian, Spanish, Portuguese) for instruction-tuned models.
Model Avg. EU21-ARC EU21-HeSw EU21-TQA EU21-MMLU
Meta-Llama-3.1-8B-Instruct .538 .529 .540 .530 .554
Mistral-7B-Instruct-v0.3 .492 .480 .485 .540 .463
Salamandra-7B-Instruct .535 .576 .619 .493 .451
Aya-23-8B .450 .420 .473 .479 .427
Occiglot-7B-eu5-Instruct .432 .419 .441 .476 .392
Pharia-1-LLM-7B-C-A .352 .296 .334 .458 .319
Bloomz-7B1 .328 .263 .295 .475 .281
Teuken-7B-Base (Ours) .511 .539 .597 .466 .443
Teuken-7B-Instruct (Ours) .559 .569 .640 .583 .443
(c) Results on multilingual benchmarks across the 15 exclusive European languages (Romanian, Czech, Danish, Greek, Estonian, Finnish, Hungarian, Lithuanian, Latvian, Dutch, Bulgarian, Polish, Slovak, Slovenian, and Swedish) for instruction-tuned models.
@@
Aya-23-8B .563 .593 .661 .484 .512
Salamandra-7B-Instruct .557 .620 .664 .476 .468
Bloomz-7B1 .394 .373 .419 .456 .327
Teuken-7B-Instruct (Ours) .585 .619 .692 .564 .464
(d) Instruct model results on multilingual benchmarks across the 10 common languages (Czech, Dutch, English, French, Greek, Italian, Polish, Portuguese, Romanian, and Spanish). The tables include mean accuracy across languages.
@@ 7.5 Comparison on Common Languages
Table˜2(d) compares the performance of Teuken-7B-Instruct (Ours) models with Aya-23-8B, Bloomz-7B1, and Salamandra-7B-Instruct across ten common European languages: Czech, Dutch, English, French, Greek, Italian, Polish, Portuguese, Romanian, and Spanish. Teuken-7B-Instruct (Ours) leads with an average of 58.5%, obtaining the best results in EU21-HeSw (69.2%) and EU21-TQA (56.4%). Aya-23-8 B’s, the second best-performing model, achieves an average performance of 56.3%, excelling on EU21-MMLU (51.2%).
Overall, the results highlight the strong performance of our model across various language sets.
@@ 7.6 Multilingual Instruction Following Evaluation
For evaluating the multilingual instruction-following capabilities of Teuken-7B-Instruct, we utilized MT-Bench-X [1] that employs an LLM as a judge across all five available evaluation languages: English, German, French, Italian, and Spanish. The results are presented in the Section˜A.5.3 and Figure˜20. We compare the cross-lingual performance of our model with Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3 and Salamandra-7B-Instruct in the following. Overall, Teuken-7B-Instruct demonstrates robust cross-lingual performance in several key domains. In particular, the Teuken-7B-Instruct model exhibits strengths in creative tasks such as Writing and Roleplay and in knowledge-based domains like Humanities and Stem, especially in German. However, it noticeably underperforms in the Math and Coding categories, highlighting areas for potential future optimization. This performance gap can be explained by the fact that our model has not been optimized for these capabilities. Compared to Salamandra-7B-Instruct, our model provides superior multilingual capabilities across most tasks and languages. Notable exceptions include the Reasoning tasks in Italian and German and the extraction task, where Salamandra-7B-Instruct achieves slightly better cross-lingual performance.
7.6.1 Toxicity
To measure the toxicity of our model, we compare our models against other instruction-tuned baselines using the PolygloToxicityPrompts benchmark [17], reporting results for the $\textsc{PTP}_{\text{SMALL}}$ subset. Our evaluation setup follows the protocol described in [17], with one exception: we set the repetition parameter to $K=1$. In addition to the benchmark’s core metrics, we report toxicity and profanity scores using the Perspective API [6]. For each attribute $i\in{\text{Profanity},\text{Toxicity}}$, we compute both the Empirical Probability ($\text{EP}_{i}$) and the Average Score ($\text{A}_{i}$). Table˜3 presents the aggregated results across five languages: German, English, French, Italian, and Spanish. Detailed language-specific scores are provided in Table˜3236 in the appendix.
Model $\textsc{EP}_{\textit{Prof.}}$ $\textsc{A}_{\textit{Prof.}}$ $\textsc{EP}_{\textit{Tox.}}$ $\textsc{A}_{\textit{Tox.}}$
Meta-Llama-3.1-8B-Instruct .174 .196 .173 .219
Mistral-7B-Instruct-v0.3 .139 .160 .140 .183
Salamandra-7B-Instruct .206 .216 .196 .230
Aya-23-8B .177 .197 .176 .218
Occiglot-7B-eu5-Instruct .202 .212 .193 .226
Bloomz-7B1 .100 .115 .106 .134
Teuken-7B-Instruct (Ours) .081 .149 .076 .152
Among the evaluated models, Teuken-7B-Instruct (Ours) demonstrates superior per- formance in minimizing the likelihood of generating harmful content. It achieves the lowest EPProfanity of 0.081 and EPToxicity of 0.076, indicating that it is the least likely to produce pro- fane or toxic responses. In comparison, Bloomz-7B1, the next best performer, has EPProfanity of 0.100 and EPToxicity of 0.106. Other models, such as Salamandra-7B-Instruct and Occiglot- 7B-eu5-Instruct, exhibit significantly higher probabilities, with EPToxicity values exceeding 0.19, The Teuken-7B-Instruct model excels in minimizing harmful content, with the lowest empirical probabilities for profanity (0.081) and toxicity (0.076) on the PolygloToxicityPrompts benchmark. This makes it less likely to generate harmful outputs compared to other models. Although Bloomz-7B1 has slightly lower severity scores when harmful content occurs (0.115 for profanity, 0.134 for toxicity) versus Teuken’s (0.149 and 0.152), Teuken’s lower occurrence rate results in a safer expected toxicity (0.0116) than Bloomz’s (0.0142). Other models, like Salamandra-7B-Instruct and Occiglot-7B-eu5-Instruct, perform worse, with expected toxicity values above 0.04. Thus, Teuken-7B-Instruct is a top choice for safety-focused applications. highlighting Teuken’s substantial improvement in this regard. In conclusion, Teuken-7B-Instruct stands out for its ability to reduce the likelihood of harmful outputs while maintaining competitive severity levels, making it a strong candidate for appli- cations prioritizing safety and reliability.
@@ 8 Conclusion & Future Work
In this work, we presented the development of two multilingual large language models, Teuken-7B-Base and Teuken-7B-Instruct, tailored to support the linguistic diversity of Europe by encompassing all 24 official EU languages. Through the use of a custom multilingual tokenizer and a dataset prioritizing non-English content, our models address limitations found in existing multilingual models, particularly their English-centric bias. Our results demonstrate strong performance across multiple benchmarks, including ARC, HellaSwag, MMLU, and TruthfulQA, while training on significantly fewer tokens compared to related models. The presented multilingual models, focusing on European languages, are an excellent base for further pre-training or fine-tuning each European language. We have taken great care to ensure that all languages are represented in the tokenizer, promoting the inclusivity and applicability of our models in diverse linguistic contexts. We shared insights regarding our model development, often not described in detail, to support future development aimed at developing multilingual models.
In the future, we aim to expand our efforts along four key dimensions. First, we plan to enhance our models’ capabilities further, focusing on mathematics and coding. To achieve this, we will continue training and fine-tuning our models, placing special emphasis on filtering high-quality reasoning, math and code data from the web—a strategy shown to be effective [60, 46]. Additionally, we will synthesize new training examples to boost performance in these domains. Second, we intend to curate even higher-quality pre-training datasets by leveraging (large) language models as judges in the filtering process, building on recent advancements [19]. Third, we aim to broaden our linguistic coverage by training on a more diverse set of European languages beyond the 24 official EU languages, making our technology accessible to an even wider audience. Finally, we plan to increase the size of the model to excel in increasingly complex tasks and push the performance boundaries.
Acknowledgments
This work was funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D) as well as by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine Learning and Artificial Intelligence, LAMARR22B, and by the European Union’s Horizon 2020 research and innovation program under grant agreement No 101135671 (TrustLLM). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) as well as the Center for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for providing its facilities for automatic evaluation computations.